HUBBLE SPACE TELESCOPE DAILY REPORT #2956 PERIOD COVERED: 0000Z (UTC) 09/13/01 - 0000Z (UTC) 09/14/01 Daily Status Report as of 257/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of WF/PC-2 9244 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. As described in HSTAR 8339 and 2.1, the acquisition for the third iteration of this proposal defaulted to fine lock backup on one FGS only, possibly affecting one observation. Otherwise, the observations completed with no further anomalous activity. 1.2 Completed Two Sets of STIS/CCD 8901 (Dark Monitor-Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. There was no anomalous activity. 1.3 Completed Four Sets of WF/PC-2 8936 (Cycle 10 Supplemental Darks Pt1/3) The WF/PC-2 was used to perform a dark calibration program that obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed with no reported problems. 1.4 Completed WF/PC-2 8584 (Calibrating the Metallicity Dependence of the Cepheid PL Relation) The WF/PC-2 was used to determine the uncertainty in the metal abundance dependence of the Cepheid PL relation that remains as one of the largest sources of systematic error in the Cepheid distance scale and in the extragalactic distance scale as a whole. There were no reported problems. 1.5 Completed STIS/CCD 9285 (POMS Test Proposal: STIS Non-scripted Parallel Proposal Continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.6 Completed STIS/CCD 9051 (Identifying Damped Lyman-alpha Galaxies at z~1) The Space Telescope Imaging Spectrograph (CCD) was used to look for samped Lyman-alpha absorption systems that contain the bulk of the neutral gas in the Universe in the redshift range z = 0.5 - 5, yet the nature of the galaxies responsible for the absorption is not well understood. Only recently have observers found more than a handful of damped absorbers at redshifts z < 1.5. Using the FIRST Bright Quasar Survey {FBQS}, with over a 1000 quasars, the proposers have undertaken a survey to build a complete picture of he nature of the galaxies responsible for damped Lyman-alpha absorption systems at z~1 and to double the sample size at this redshift. No problems were encountered. 1.7 Completed STIS/CCD 8903 (Bias Monitor - Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 to build up high-S/N superbiases and track the evolution of hot columns. There were no problems. 1.8 Completed WF/PC-2/STIS/CCD 9174 (Using Optically Faint Radio Sources to Pinpoint Dusty Proto-Galaxies) The WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to observe a number of high redshift, dusty starburst galaxies that remain invisible in ground based in the Hubble Deep Field. These galaxies appear as faint radio sources with bright submillimeter counterparts, demonstrating they are very dusty systems. The luminosities of these galaxies suggest they are in the process of converting the bulk of their gas mass into stars, and will likely evolve into present day massive ellipticals. The observations completed normally, with no reported problems. 1.9 Completed STIS/CCD 9136 (T Tauri Star Coronagraphic Survey: A PMS Protoplanetary Disk Census) The Space Telescope Imaging Spectrograph (CCD) was used to observe pre-main sequence solar-mass stars, the T Tauri stars that Millimeter and IR studies suggest at least 50 percent have circumstellar disks similar to the disk from which our planetary system formed. High spatial resolution, high dynamic range imaging of such systems will map the spatial distribution of material around the star, constraining the disk sizes and inclinations, and provide a first assessment of when structure in the disk, such as cleared central zones and annuli, which has been linked to planet formation, develops. As described in HSTAR 8339 and 2.1, the acquisition for this proposal defaulted to fine lock backup on one FGS only, possibly affecting five observations. Otherwise, all observations completed without incident. 1.10 Completed WF/PC-2 8941 (Cycle 10 UV Earthflats) The WF/PC-2 was used to monitor flat field stability by obtaining sequences of earth streak flats to improve the quality of pipeline flat fields for the WFPC2 UV filter set. As described in HSTAR 8339 and 2.1, the acquisition for this proposal defaulted to fine lock backup on one FGS only, possibly affecting two observations. Otherwise, the proposal had no other problems. 1.11 Completed Three Sets of STIS/CCD 9088 (Next Generation Spectral Library of Stars) The Space Telescope Imaging Spectrograph (CCD) was used to produce a "Next Generation'' Spectral Library of 600 stars for use in modeling the integrated light of galaxies and clusters by using the low dispersion UV and optical gratings of STIS. The library will be roughly equally divided among four metallicities, very low {Fe/H < -1.5}, low {-1.5 < Fe/H < -0.5}, near-solar {-0.5 < Fe/H < 0.1}, and super-solar {Fe/H > 0.1}, well-sampling the entire HR-diagram in each bin. Such a library will surpass all extant compilations and have lasting archival value, well into the Next Generation Space Telescope era. No problems occurred. 1.12 Completed STIS/CCD 8911 (Sparse-Field CTE External C10) The Space Telescope Imaging Spectrograph (CCD) was used to determine the derivation of accurate coefficients which can be used to correct low count level data {both imaging and spectroscopy} for nonlinearity/CTE effects. Dependences on x-position, background sky level, and time dependence will be quantified in addition to accurate determination of intensity and y-position effects in a third epoch and to higher accuracy than previously determined. The proposal completed with no reported problems. 1.13 Completed Three Sets of WF/PC-2 9249 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to execute a POMS test proposal designed to simulate scientific plans. The proposal completed with no reported problems. 1.14 Completed WF/PC-2 9160 (Disks and Envelopes of Nearby Nebulous Young Stellar Objects: A Snapshot Survey) The WF/PC-2 was used to perform a snapshot survey of nearby nebulous young stellar objects to study the detailed morphology of the their disks and envelopes and probe the effect of inclination on the infrared spectral energy distribution of disk/envelope systems. The proposal completed as planned. 1.15 Completed STIS/CCD/MA2 9040 (Baryons In Intermediate Redshift {Z > 1} OVI Absorbers) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to search for intervening OVI absorption systems in two further extremely UV bright intermediate redshift QSOs {HS 0747 + 4259, z = 1.9, V = 15.6; HS 0818+2554, z = 1.5, V = 15.4} using the STIS E230M Echelle mode. The scientific aim is to measure the baryonic fraction in the warm-hot intergalactic medium at redshifts 1 <=q z <=q 1.8. Present existing data on HE 0515-4414 taken also with STIS E230M seem to indicate that the baryonic fraction at z = 1.5 is lower by a factor of ~ 15 than the high value found by Tripp et al. {2000} for z < 0.3. Our results need to be confirmed by more lines of sight, and we plan to test hierarchical structure formation models which predict a rapid evolution between z = 1.5 and z = 0. The observations completed as planned with no reported anomalous activity. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 8 Successful: 8 Per HSTAR 8339, the acquisition at 256/115341Z defaulted to fine lock backup on FGS-1 when the scan step limit was exceeded on FGS-3. The science proposals detailed in 1.1, 1.9 and 1.10 may have been affected. Scheduled Re-acquisitions: 6 Successful: 6 2.2 FHST Updates: Scheduled: 17 Successful: 17 2.3 Operations Notes: The SSR EDAC error counter was cleared twice, per ROP SR-1A. A TTR Was generated when there was an engineering data loss from 256/1009Z until 256/1029Z due to bad weather. An SSR engineering record was performed at 256/1032Z per ROP SR-4A. The NSSC-1 status buffer was dumped and reset at 257/0507Z as directed by ROP NS-3. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW